# Doping $S=1$ antiferromagnet in one-dimension

### v1

@@ IV.2.2 Interaction dependence
Let us finally comment on the interaction $U$ dependence of spin excitations for finite doping $x$. The results for half-filling ($x=0$) were presented in Ref. [43]. Here, in Fig. 7, we focus on $x=0.2$.
@@
At small interaction (below the orbital-RVB phase), the spin excitations resemble those of the single-orbital Hubbard model, i.e., the two-spinon continuum [see Fig. 7(a) for $U/W=1$ result]. Such behavior is expected, as in the $U\to 0$ limit the two-orbital HK model reduces to two noninteracting chains. In the latter case, the boundaries of the spin spectrum can be evaluated using the Lindhard formula [63]. Increasing the interaction strength in the single-orbital Hubbard model leads to $S=1/2$ Heisenberg [64] or $t$-$J$ model [31] at $x=0$ or $x\neq 0$, respectively. In the former case, a two-spinon continuum can be calculated with the help of Bethe Ansatz [65]. Both solutions, $U\to 0$ and $U\to\infty$, have the same functional form [e.g., for $x=0$ and zero magnetization: $\omega_{\mathrm{U}}(k)=\alpha|\sin(k/2)|$ and $\omega_{\mathrm{L}}(k)=\alpha/2|\sin(k)|$], but with different energy scales $\alpha$, i.e., $\alpha=4t$ and $\alpha=\pi 4t^{2}/U$, respectively. Note also that the latter value ($U\to\infty$ prediction) is not valid for finite $U$, especially for $U\sim W$ [64]. Furthermore, the two limits have different distributions of spectral weights with a maximum of the intensity at the top (bottom) of the continuum in the $U\to 0$ ($U\to\infty$) limit, and as a consequence, increasing the interaction strength leads to a gradual transfer part of weight to the $\omega_{\mathrm{L}}$ line, preserving overall continuum structure [31, 64].
@@
Figure 7: Interaction $U$ dependence of the dynamical spin structure factor $S(q,\omega)$ of the two-orbital Hubbard model for $x=0.2$. Panels (a-d) depict data for $U/W=1,2,3,4$, respectively. The white dashed line in panel (a) depicts the upper and lower bounds of the two-spinon continuum of the $t$-$J$ model at finite doping. In each plot, 100100 frequency points are shown ($\Delta\omega=\omega_{\mathrm{max}}/100$, where $\omega_{\mathrm{max}}$ is the maximal presented frequency). Left $y$-axis represents frequency $\omega$ in unit of hopping $t$, while right $y$-axis represents $\omega$ in units of $J=2t^{2}/(U+J_{\mathrm{H}})$.
@@
Finally, an interesting aspect of our findings is that further increasing the interaction strength has a drastic effect on the spin excitations. Close to the transition between the orbital-RVB and ferromagnetic phase [$U/W=4$, Fig. 7(d)], i.e., in the region of parameters where the maximum $q_{\mathrm{max}}$ of static SSF $S(q)$ does not follow the Fermi vector prediction, $q_{\mathrm{max}}<2k_{\mathrm{F}}$ [see Fig. 4(d)]. Here, the spin spectrum resembles that of the $J_{1}$-$J_{2}$ Heisenberg model with AFM couplings $J_{1}\simeq J_{2}$ [66]. The latter can give support to the spiral phase, in agreement with our static SSF investigation presented in Sec. IV.1. Note that a similar spiral phase was reported in the OSMP system, also as a precursor of the ferromagnetic phase [59, 61].
@@ V Conclusions
We have analyzed the ground state and the static $S(q)$ and dynamic $S(q,\omega)$ spin structure factor, i.e., the spin ordering and spin excitations, in the hole-doped $S=1$ Heisenberg model. The latter have to be investigated using the two-orbital Hubbard-Kanamori Hamiltonian, as opposed to the $S=1/2$ case, where the single-orbital Hubbard model and $t$-$J$ model capture the overall physics. Primarily, we have focused on the so-called orbital-RVB phase, i.e., the valence-bond-like ground state at finite doping with properties inherited from the AKLT state of the $S=1$ Heisenberg model.
@@
From an experimental perspective, our predictions could be directly tested in quasi-one-dimensional $S=1$ compounds such as doped nickelate chains (e.g., vacancy- or hole-doped Y2-xCaxBaNiO5 [17, 26, 27]) using INS, where both the doping-dependent shift of the magnon maximum from $q=\pi$ to $q=2k_{\mathrm{F}}$ and the emergence of high-frequency spectral weight for $q\gtrsim 2k_{\mathrm{F}}$ should be observable. Resonant inelastic X-ray scattering (RIXS) [23, 29, 24], could provide complementary access to the short-wavelength regime, enabling detection of the broad incoherent modes analogous to those reported in orbital-selective Mott compounds such as BaFe2Se3 [67, 57]. Moreover, cold-atom quantum simulators of multiorbital Hubbard chains [12, 13] offer a clean and tunable platform to probe the orbital-RVB regime directly, allowing for systematic variation of interaction strength and doping to map the phase boundaries discussed here. Such combined experimental efforts would provide stringent tests of the coexistence of topological properties, coherent magnons, and incoherent high-energy excitations predicted in this work.

### v2

@@ IV.2.2 Hubbard interaction dependence
In this section, comment on the interaction $U$ dependence of spin excitations for finite doping $x$. The results for half-filling ($x=0$) were presented in Ref. [45]. Here, in Fig. 7, we focus on $x=0.2$.
@@
At small interaction (below the orbital-RVB phase), the spin excitations resemble those of the single-orbital Hubbard model, i.e., the two-spinon continuum [see Fig. 7(a) for $U/W=1$ result]. Such behavior is expected, as in the $U\to 0$ limit the two-orbital HK model reduces to two noninteracting chains. In the latter case, the boundaries of the spin spectrum can be evaluated using the Lindhard formula [66]. In the case of single-orbital Hubbard mode, increasing the interaction strength leads to $S=1/2$ Heisenberg [67] or $t$-$J$ model [31] at $x=0$ or $x\neq 0$, respectively. In the former case, a two-spinon continuum can be calculated with the help of the Bethe Ansatz [68]. Both solutions, $U\to 0$ and $U\to\infty$, have the same functional form [e.g., for $x=0$ and zero magnetization: $\omega_{\mathrm{U}}(k)=\alpha|\sin(k/2)|$ and $\omega_{\mathrm{L}}(k)=\alpha/2|\sin(k)|$], but with different energy scales $\alpha$, i.e., $\alpha=4t$ and $\alpha=\pi 4t^{2}/U$, respectively. Note also that the latter value ($U\to\infty$ prediction) is not valid for finite $U$, especially for $U\sim W$ [67]. Furthermore, the two limits have different distributions of spectral weights with a maximum of the intensity at the top (bottom) of the continuum in the $U\to 0$ ($U\to\infty$) limit, and as a consequence, increasing the interaction strength leads to a gradual transfer part of weight to the $\omega_{\mathrm{L}}$ line, preserving overall continuum structure [31, 67].
@@
Figure 7: Interaction $U$ dependence of the dynamical spin structure factor $S(q,\omega)$ for finite doping $x=0.2$. Panels (a-e) depict data for $U/W=1,2,3,4,8$, respectively. The white dashed line in panel (a) depicts the upper and lower bounds of the two-spinon continuum of the $t$-$J$ model at finite doping. The white dashed line in panel (e) depicts the standard magnon dispersion $\omega\propto 1-\cos(q)$ above ferromagnetically order state. In each plot, 100100 frequency points are shown ($\Delta\omega=\omega_{\mathrm{max}}/100$, where $\omega_{\mathrm{max}}$ is the maximal presented frequency). Left $y$-axis represents frequency $\omega$ in unit of hopping $t$, while right $y$-axis represents $\omega$ in units of $J=2t^{2}/(U+J_{\mathrm{H}})$.
@@
An interesting aspect of our findings is that further increasing the interaction strength has a drastic effect on the spin excitations. Close to the transition between the orbital-RVB and ferromagnetic phase [$U/W=4$, Fig. 7(d)], i.e., in the region of parameters where the maximum $q_{\mathrm{max}}$ of static SSF $S(q)$ does not follow the Fermi vector prediction, $q_{\mathrm{max}}<2k_{\mathrm{F}}$ [see Fig. 4(d)]. Here, the spin spectrum resembles that of the $J_{1}$-$J_{2}$ Heisenberg model with AFM couplings $J_{1}\simeq J_{2}$ [69]. The latter can give support to the spiral phase, in agreement with our static SSF investigation presented in Sec. IV.1. Note that a similar spiral phase was reported in the OSMP system, also as a precursor of the ferromagnetic phase [61, 63].
Finally, for completeness, in Fig. 7(e) we present results for the ferromagnetic phase (generically placed at $U\gg t$ and $x\neq 0$; here considered at $U/W=8$ and $x=0.2$). In such a kinematically stabilized phase, electrons on orbital $\gamma$ minimize their kinetic energy when they can hop on the site where the electron on orbital $\gamma\prime$ has the same spin projection - thus leading to an overall ferromagnetic spin arrangement. Previous investigation of the systems with lifted orbital degeneracy (e.g., within OSMP phase) has shown [55] that the spin excitations above double-exchange ferromagnets follow the ”standard” (Holstein–Primakoff-like) magnon dispersion, $\omega\propto 1-\cos(q)$, only in the limit of long-wavelengths $q\to 0$. For short-wavelengths $q>\pi/2$, one observes a broad incoherent excitations due to scattering on the Stoner continuum. Our results presented in Fig. 7(e) indicate that the same behavior can be observed in the case of degenerated bands at $U\gg t$. It is worth noting that the overall energy scale of the spin excitation in this phase does not depend strongly on the values of the Hubbard or Hund interaction, but rather on the kinetic energy of the electrons (controlled by, i.e., by the density of the electrons $n$).
IV.2.3 Hund interaction dependence
Since the overall energy scale of the spin excitations in orbital-RVB phase is given by $J=2t^{2}/(U+J_{\mathrm{H}})$, the value of the Hund interaction ($J_{\mathrm{H}}/U$ ratio in our notation) can also influence the overall behavior. For example, one can expect that for a large Hubbard interaction $U$ but vanishing Hund exchange $J_{\mathrm{H}}\to 0$ the spin excitations $S(q,\omega)$ should be akin to a single-band $t$-$J$ result (i.e., two-spinon continuum). Our results for $J_{\mathrm{H}}/U=0.1$, presented in Fig. 8(a), indeed confirm such scenario. The latter is also expected from the analysis of the $J_{\mathrm{H}}$-$U$ phase diagram [45]. I.e., for $x=0$, the transition between single-band $S=1/2$ and two-orbital $S=1$ physics is given by $J_{\mathrm{H}}=t^{2}/U$. Since for finite doping $x\neq 0$ the transition occurs [40] at a larger value of interaction $U$, and can expect that a similar relation $J_{\mathrm{H}}\propto b\,t^{2}/U$, with $b>1$.
For sufficiently large values of the Hund interaction, the system enters the orbital-RVB phase. Our results indicate that within this phase, the results don’t depend strongly on the specific value of $J_{\mathrm{H}}/U$. Such a behaviour is presented in Fig. 8(b,c), where we present $S(q,\omega)$ for $J_{\mathrm{H}}/U=0.25$ and $0.4$ respectively. Our results indicate that the low-frequency (magnon-like) part is just renormalized by the change in the effective spin exchange value $J=J(J_{\mathrm{H}})$, as expected. Similarly, the high-frequency incoherent weight at $q>2k_{\mathrm{F}}$ sharpens, but overall ”continuum” is still clearly visible.
Figure 8: Hund interaction $J_{\mathrm{H}}/U$ dependence of the dynamical spin structure factor $S(q,\omega)$ for finite doping $x=0.4$. Panels (a-c) depict results for $J_{\mathrm{H}}/U=0.10,0.25,0.40$, respectively. The results below (above) the white dashed line are evaluated with frequency resolution $\Delta\omega/t=0.04$ ($\Delta\omega/t=0.004$).
It is important to note that for sufficiently large of Hubbard interaction $U$ (and consequently for large $J_{H}$) the system will enter the ferromagnetic phase [55]. For the values considered in this work (e.g., for $U/W=3$ presented in Fig. 8) the system remains in the orbital-RVB phase up to maximal possible $J_{\mathrm{H}}/U=0.4$ (for $J_{\mathrm{H}}/U>0.4$ the interorbital interaction $U^{\prime}$ takes a nonphysical negative value). However, for elevated values of interaction $U\gg t$, we expect that the orbital-RVB phase is stable only in a narrow (non-zero) range of $J_{\mathrm{H}}/U$ ratio, giving way to the ferromagnetic phase for larger $J_{\mathrm{H}}$.
@@ V Conclusions
We have analyzed the ground state and the static $S(q)$ and dynamic $S(q,\omega)$ spin structure factor, i.e., the spin ordering and spin excitations, in the hole-doped $S=1$ Heisenberg model. The latter have to be investigated using the two-orbital Hubbard-Kanamori Hamiltonian, rather than the $S=1/2$ case, where the single-orbital Hubbard model and the $t$-$J$ model capture the overall physics. Primarily, we have focused on the so-called orbital-RVB phase, i.e., the valence-bond-like ground state at finite doping with properties inherited from the AKLT state of the $S=1$ Heisenberg model.
@@
Although experimentally controlled carrier doping in $S=1$ Haldane chains is challenging, several material platforms demonstrate that this regime is physically meaningful and experimentally accessible [17, 26, 27]. In particular, hole-doped nickelate chains such as Y2-xSrxBaNiO5 have long been known to realize doped $S=1$ backgrounds; photoemission spectroscopy shows a significant increase in spectral weight at the top of the valence band upon doping [70]. Early theoretical work [71] on the nickelate chain compound Y2BaNiO5 demonstrated that a multiband model including oxygen and nickel orbitals gives rise to mobile holes interacting with a $S=1$ spin chain, producing sub-gap states in the dynamical spin structure factor consistent with experimental observations. Furthermore, magnetization and susceptibility studies [72] on Ca-doped Y2-xCaxBaNiO5 reveal a crossover to a spin-glass–like state at low temperatures, confirming that doping strongly modifies magnetic correlations in the $S=1$ chain background. Consequently, while fully tunable doping in ideal $S=1$ chains remains experimentally rare, existing solid-state nickelate compounds already enter regimes where predictions such as doping-dependent shifts of magnon spectral weight and incommensurate magnetic correlations should be testable. Beyond solid-state materials, cold-atom quantum simulators of multiorbital Hubbard systems have recently achieved the Haldane phase [12, 13] and allow tunable doping through particle imbalance. Thus, while fully tunable doping in ideal $S=1$ chains is experimentally rare, several platforms already enter regimes where the physics uncovered here - such as the persistence of a gapped magnon branch, the doping-dependent shift of the spectral maximum to $q=2k_{\mathrm{F}}$, and the emergence of broad high-energy spectral weight - should be detectable.
The behavior we uncover also has potential implications for higher-dimensional doped antiferromagnets, particularly multiorbital systems where $S>1/2$ local moments arise from Hund’s coupling [21, 22, 44]. A central question in two-dimensional correlated materials - including iron-based superconductors and doped nickelates - is how itinerant carriers propagate in an antiferromagnetic background and whether coherent magnon excitations survive in the presence of doping. The latter question are encapsulated in the numerus studies of doped (two-dimensional) spin liquids [39, 73]. Our results demonstrate that (at least in 1D), ”coherent” long-wavelength magnons and ”incoherent” spinons coexist within an orbital-RVB phase. This coexistence mirrors aspects of the “Hund-metal” phenomenology [25] in higher dimensions, where local-moment physics and itinerancy intertwine.